OpenStudy (anonymous):

multiply (x^3/8)(x^2/5)

OpenStudy (anonymous):

you have options?

OpenStudy (anonymous):

its x^5/40

OpenStudy (anonymous):

a. x^1/4 b. x^5/13 c. x^3/20 d. x^31/40

OpenStudy (sedatefrog712):

I think it is D but Im not 100% sure

OpenStudy (anonymous):

i got D initially but i wasn't sure either...

OpenStudy (anonymous):

I would say D

OpenStudy (sedatefrog712):

well id say that it is D

OpenStudy (anonymous):

\[x ^{\frac{ 3 }{ 8 }} × x ^{\frac{ 2 }{ 5 }}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply 3/8 by 5 to get the LCM for denominator so that exponent is 15/40 then multiply 2/5 by 8 to get the lcm 16/40 then add exponents \[x ^{\frac{ 31 }{ 40 }}\]

OpenStudy (sedatefrog712):

yep

OpenStudy (anonymous):

thank you1

OpenStudy (anonymous):

you're welcome

OpenStudy (sedatefrog712):

Welcome im always happy to help